# Tracial states on groupoid $C^{*}$-algebras and essential freeness

### v2

@@ 4. Extension of $\tau^{\mathrm{Fix}}_{\mu}$
We recall the following special case of [26, Theorem 1.1 and Theorem 1.3]:
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second countable, and $\tau$ be a tracial state on $C^{*}(\mathcal{G})$ with the associated measure $\mu$. Then there exists a unique $\mu$-measurable field $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$, where $\tau_{x}$ is a tracial state on $C^{*}(\mathcal{G}_{x}^{x})$ such that
[…]
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid, and $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. For each $x\in\mathcal{G}^{(0)}$, we denote $\tau_{x}^{\mathrm{triv}}:C_{c}(\mathcal{G}_{x}^{x})\to\mathbb{C}$ the trivial representation of $\mathcal{G}_{x}^{x}$. Then for every $f\in C_{c}(\mathcal{G})$, we have
@@
It suffices to verify (4.2) for $g\in C_{c}(B)$, where $B$ is a pre-compact open bisection. By definition, $\tau^{\mathrm{triv}}_{x}(\eta_{x}(g))\neq 0$ if and only if there exists $\gamma\in B$ such that $\mathrm{s}(\gamma)=\mathrm{r}(\gamma)$ and $g(\gamma)\neq 0$. This happens if and only if $\mathrm{s}(\gamma)\in\mathrm{Fix}(\alpha_{B})$ and $g(\gamma)\neq 0$. Hence, we obtain
@@
which concludes the proof. ∎
@@
It is worth noticing that we do not need second countability of $\mathcal{G}$ in the hypothesis of Lemma 4.2. The reason is that we can simply find the associated measurable field of tracial states $\{\tau_{x}^{\mathrm{triv}}\}_{x\in\mathcal{G}^{(0)}}$ and then directly verify (4.2) without using Renault’s disintegration theorem. However, Renault’s disintegration theorem was used in the proof of Proposition 4.1.
@@
Sergey Neshveyev has pointed out to us that the fixed point trace $\tau^{\mathrm{Fix}}_{\mu}$ in the current paper coincides with $\varphi_{\mu}^{\prime\prime}$ considered in [27, Equation (2.3)] by Lemma 4.2.
@@
Denote $C^{*}_{e}(\mathcal{G}_{x}^{x})$ the $C^{*}$-completion of $C_{c}(\mathcal{G}^{x}_{x})$ with respect to $\|\cdot\|_{e}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second countable, and $\tau$ be a tracial state on $C^{*}(\mathcal{G})$ with the associated measure $\mu$ on $\mathcal{G}^{(0)}$. Then $\tau$ factors through $C^{*}_{r}(\mathcal{G})$ if and only if for the associated $\mu$-measurable field of tracial states $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$ (as in Proposition 4.1), $\tau_{x}$ factors through $C^{*}_{e}(\mathcal{G}_{x}^{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second countable, and $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. Then $\tau^{\mathrm{Fix}}_{\mu}$ extends to a tracial state on $C^{*}_{r}(\mathcal{G})$ if and only if the trivial representation $\tau_{x}^{\mathrm{triv}}$ on $C_{c}(\mathcal{G}_{x}^{x})$ extends to a tracial state on $C^{*}_{e}(\mathcal{G}_{x}^{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second countable, and $\tau$ be a tracial state on $C^{*}_{r}(\mathcal{G})$ with the associated measure $\mu$ on $\mathcal{G}^{(0)}$. If $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}(\mathcal{G}^{x}_{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$, then for the associated $\mu$-measurable field of tracial states $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$ (as in Proposition 4.1), $\tau_{x}$ factors through a state $\tau_{x,r}$ on $C^{*}_{r}(\mathcal{G}_{x}^{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second countable, and $\tau$ be a tracial state on $C^{*}_{r}(\mathcal{G})$ with the associated measure $\mu$ on $\mathcal{G}^{(0)}$. Assume that for $\mu$-almost every $x\in\mathcal{G}^{(0)}$, we have $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}(\mathcal{G}^{x}_{x})$. Then for $a\in C^{*}_{r}(\mathcal{G})$, we have
@@
for sufficiently large $n$. Hence, Lebesgue’s dominating convergence theorem implies that
@@
Using an identical argument as in the proof of Proposition 4.8, (4.1) can be always extended from $C_{c}(\mathcal{G})$ to $C^{*}(\mathcal{G})$ without requiring $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}(\mathcal{G}^{x}_{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$. More precisely,
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second countable, and $\tau$ be a tracial state on $C^{*}(\mathcal{G})$ with the associated measure $\mu$. Then there exists a unique $\mu$-measurable field $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$, where $\tau_{x}$ is a tracial state on $C^{*}(\mathcal{G}_{x}^{x})$ such that
[…]
@@ 5.1. Transformation groupoids
In this subsection, we apply the results in previous sections to transformation groupoids. If $\Gamma$ is a discrete group acting on a locally compact Hausdorff space $X$, then we form the transformation groupoid $X\rtimes\Gamma$. In this case, the reduced groupoid $C^{*}$-algebra $C^{*}_{r}(X\rtimes\Gamma)$ is ***-isomorphic to the reduced crossed product $C_{0}(X)\rtimes_{r}\Gamma$, and the maximal groupoid $C^{*}$-algebra $C^{*}(X\rtimes\Gamma)$ is ***-isomorphic to the maximal crossed product $C_{0}(X)\rtimes\Gamma$.
For each $\gamma\in\Gamma$, we use the same notation to denote the associated homeomorphism $\gamma:X\to X$. Therefore, a probability Radon measure $\mu$ on $X$ is invariant in the sense in Section 2.3 if and only if $\gamma$ is $\mu$-preserving for each $\gamma\in\Gamma$. Applying the inner regularity of Radon measures, the notion of essential freeness for transformation groupoids can be easily simplified as follows:
@@
Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ with an invariant probability Radon measure $\mu$ on $X$. Then the transformation groupoid $X\rtimes\Gamma$ is essentially free with respect to $\mu$ (in the sense of Definition 3.3) if and only if $\mu(\mathrm{Fix}(\gamma))=0$ for every $\gamma\neq 1$ in $\Gamma$.
@@
where $f_{i}\in C_{0}(X)$ and $\gamma_{i}\in\Gamma$ for $i=1,\cdots,n$.
@@
Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ equipped with an invariant probability Radon measure $\mu$. Consider the following conditions:
@@
Then (2) normal-$\Rightarrow$ (1).
If $\Gamma$ is countable and $X$ is second countable, then we actually have (1) normal-$\Leftrightarrow$ (2).
@@
Firstly, we show that “(2) $\Rightarrow$ (1)”. Recall from Lemma 4.2 (see also Remark 4.3) that
@@
holds for any $f\in C_{c}(\mathcal{G})$. By (2), the trivial representation $\tau^{\mathrm{triv}}_{x}$ on $C_{c}(\Gamma_{x})$ extends to a state $\tau^{\mathrm{triv}}_{x,r}$ on $C^{*}_{r}(\Gamma_{x})$ for $\mu$-almost every $x\in X$. By the same argument as in the proof of Proposition 4.8 without requiring the second countability of $X\rtimes\Gamma$, the map
@@
Now we show that “(1) $\Rightarrow$ (2)” if $\Gamma$ is countable and $X$ is second countable. By Proposition 4.6, $\tau^{\mathrm{Fix}}_{\mu}$ extends to a tracial state on $C_{0}(X)\rtimes_{r}\Gamma$ if and only if for $\mu$-almost every $x\in X$, the trivial representation on $C_{c}((X\rtimes\Gamma)_{x}^{x})$ extends to a state on $C^{*}_{e}((X\rtimes\Gamma)_{x}^{x})$. Note that $(X\rtimes\Gamma)_{x}^{x}$ is isomorphic to $\Gamma_{x}$, and it follows from [11, Proposition 2.10] that $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}((X\rtimes\Gamma)_{x}^{x})$. It is also known that the trivial representation on $C_{c}(\Gamma_{x})$ extends to a state on $C^{*}_{r}(\Gamma_{x})$ if and only if $\Gamma_{x}$ is an amenable group. Thus, we have concluded the proof. ∎
@@
Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ equipped with an invariant probability Radon measure $\mu$. We consider the following conditions:
@@
Then (1) normal-$\Rightarrow$ (2) and (2) + (3) normal-$\Rightarrow$ (1). In particular, if the action is essentially free then any tracial state $\tau$ on $C_{0}(X)\rtimes_{r}\Gamma$ is canonical.
If additionally $\Gamma$ is countable and $X$ is second countable, then (1) normal-$\Leftrightarrow$ (2) + (3).
@@
Note that “(2) $\Rightarrow$ (1)” in Corollary 5.5 does not hold in general. Indeed, it was shown in [8, Corollary 1.4] that if the amenable radical of $\Gamma$ (i.e., the largest amenable normal subgroup of $\Gamma$) is trivial (e.g., $\Gamma$ is $C^{*}$-simple), then any tracial state on the reduced crossed product is canonical. Actually, it was shown in [33, Corollary 1.12] that any tracial state on the reduced crossed product is canonical if and only if the action of the amenable radical of $\Gamma$ on $X$ is essentially free.
@@
Let $\Gamma$ be a discrete group acting trivially on a compact Hausdorff space $X$ equipped with a probability Radon measure $\mu$. Then the following are equivalent:
@@
“(2) $\Leftrightarrow$ (3)”: If $\Gamma$ is amenable, then $X\rtimes\Gamma$ is amenable and hence has the weak containment property. Conversely, we assume that $X\rtimes\Gamma$ has the weak containment property, then the canonical quotient map $C(X)\otimes_{\text{max}}C^{*}(\Gamma)=C(X)\rtimes\Gamma\to C(X)\rtimes_{r}\Gamma=C(X)\otimes_{\text{min}}C_{r}^{*}(\Gamma)$ is injective, which implies that $C^{*}(\Gamma)\to C^{*}_{r}(\Gamma)$ is injective by a diagram chase. Therefore, $\Gamma$ is amenable. ∎
@@
Let $\Gamma$ be a non-trivial discrete group acting trivially on a compact Hausdorff space $X$ equipped with a probability Radon measure $\mu$. Then at least one of the following conditions fails:

### v3

@@ 4. Extension of $\tau^{\mathrm{Fix}}_{\mu}$
We recall the following special case of [27, Theorem 1.1 and Theorem 1.3]:
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\tau$ be a tracial state on $C^{*}(\mathcal{G})$ with the associated measure $\mu$. Then there exists a unique $\mu$-measurable field $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$, where $\tau_{x}$ is a tracial state on $C^{*}(\mathcal{G}_{x}^{x})$ such that
(4.1) $\tau(f)=\int_{\mathcal{G}^{(0)}}\sum_{\gamma\in\mathcal{G}_{x}^{x}}f(\gamma)\tau_{x}(u_{\gamma})\mathrm{d}\mu(x)\quad\text{for}\quad f\in C_{c}(\mathcal{G}).$
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid, and $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. For each $x\in\mathcal{G}^{(0)}$, we denote by $\tau_{x}^{\mathrm{triv}}:C_{c}(\mathcal{G}_{x}^{x})\to\mathbb{C}$ the trivial representation of $\mathcal{G}_{x}^{x}$. Then for every $f\in C_{c}(\mathcal{G})$, we have
@@
It suffices to verify (4.2) for $g\in C_{c}(B)$, where $B$ is a pre-compact open bisection. By definition, $\tau^{\mathrm{triv}}_{x}(\eta_{x}(g))\neq 0$ if and only if there exists $\gamma\in B$ such that $\mathrm{s}(\gamma)=\mathrm{r}(\gamma)=x$ and $g(\gamma)\neq 0$. This happens if and only if $\mathrm{s}(\gamma)\in\mathrm{Fix}(\alpha_{B})$ and $g(\gamma)\neq 0$. Hence, we obtain
@@
which concludes the proof of (4.2). The argument also shows that for such $g\in C_{c}(B)$, the map $x\mapsto\tau^{\mathrm{triv}}_{x}(\eta_{x}(g))=\chi_{\mathrm{Fix}(\alpha_{B})}(x)\cdot g((s|_{B})^{-1}(x))$ is $\mu$-measurable. Hence, the family $\{\tau^{\mathrm{triv}}_{x}\}_{x\in\mathcal{G}^{(0)}}$ is $\mu$-measurable. ∎
@@
It is worth noticing that we do not need second-countability of $\mathcal{G}$ in the hypothesis of Lemma 4.2. The reason is that we can simply find the associated measurable field of tracial states $\{\tau_{x}^{\mathrm{triv}}\}_{x\in\mathcal{G}^{(0)}}$ and then directly verify (4.2) without using Renault’s disintegration theorem. However, Renault’s disintegration theorem was used in the proof of Proposition 4.1.
@@
S. Neshveyev has pointed out to us that the fixed point trace $\tau^{\mathrm{Fix}}_{\mu}$ in the current paper coincides with $\varphi_{\mu}^{\prime\prime}$ considered in [28, Equation (2.3)] by Lemma 4.2.
@@
Denote by $C^{*}_{e}(\mathcal{G}_{x}^{x})$ the $C^{*}$-completion of $C_{c}(\mathcal{G}^{x}_{x})$ with respect to $\|\cdot\|_{e}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\tau$ be a tracial state on $C^{*}(\mathcal{G})$ with the associated measure $\mu$ on $\mathcal{G}^{(0)}$. Then $\tau$ factors through $C^{*}_{r}(\mathcal{G})$ if and only if for the associated $\mu$-measurable field of tracial states $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$ (as in Proposition 4.1), $\tau_{x}$ factors through $C^{*}_{e}(\mathcal{G}_{x}^{x})$ (denoted by $\tau_{x,e}$) for $\mu$-almost every $x\in\mathcal{G}^{(0)}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. Then $\tau^{\mathrm{Fix}}_{\mu}$ extends to a tracial state on $C^{*}_{r}(\mathcal{G})$ if and only if the trivial representation $\tau_{x}^{\mathrm{triv}}$ on $C_{c}(\mathcal{G}_{x}^{x})$ extends to a tracial state on $C^{*}_{e}(\mathcal{G}_{x}^{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$.
Consequently, we have the following:
Corollary 4.7.
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\mu$ be an invariant probability Rodan measure on $\mathcal{G}^{(0)}$. Then $\mathcal{G}$ is essentially free with respect to $\mu$ if and only if the following conditions hold:
  • $\mathcal{G}^{x}_{x}$ is amenable for $\mu$-almost all $x\in\mathcal{G}^{(0)}$;
  • for any tracial state $\tau$ on $C^{*}_{r}(\mathcal{G})$ with associated measure $\mu$, $\tau_{x,e}$ is faithful on $C^{*}_{e}(\mathcal{G}^{x}_{x})$ for $\mu$-almost all $x\in\mathcal{G}^{(0)}$.
Proof.
By definition, $\mathcal{G}$ is essentially free with respect to $\mu$ if and only if $\mathcal{G}_{x}^{x}$ is trivial for $\mu$-almost all $x\in\mathcal{G}^{(0)}$, which concludes the proof of the necessity. For sufficiency, we consider the fixed point trace $\tau_{\mu}^{\mathrm{Fix}}$. By Proposition 4.6 and the assumption of amenability of $\mathcal{G}^{x}_{x}$ for $\mu$-almost all $x\in\mathcal{G}^{(0)}$, $\tau_{\mu}^{\mathrm{Fix}}$ extends to a tracial state on $C^{*}_{r}(\mathcal{G})$. Moreover for $x\in\mathcal{G}^{(0)}$, $(\tau_{\mu}^{\mathrm{Fix}})_{x,e}$ is the trivial representation by Lemma 4.2. Hence, it is faithful if and only if it is injective, which is equivalent to that $\mathcal{G}_{x}^{x}$ is trivial. ∎
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\tau$ be a tracial state on $C^{*}_{r}(\mathcal{G})$ with the associated measure $\mu$ on $\mathcal{G}^{(0)}$. If $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}(\mathcal{G}^{x}_{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$, then for the associated $\mu$-measurable field of tracial states $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$ (as in Proposition 4.1), $\tau_{x}$ factors through a state $\tau_{x,r}$ on $C^{*}_{r}(\mathcal{G}_{x}^{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$.
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\tau$ be a tracial state on $C^{*}_{r}(\mathcal{G})$ with the associated measure $\mu$ on $\mathcal{G}^{(0)}$. Assume that for $\mu$-almost every $x\in\mathcal{G}^{(0)}$, we have $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}(\mathcal{G}^{x}_{x})$. Then for $a\in C^{*}_{r}(\mathcal{G})$, we have
@@
for sufficiently large $n$. Hence, Lebesgue’s dominated convergence theorem implies that
@@
Using an identical argument as in the proof of Proposition 4.9, (4.1) can be always extended from $C_{c}(\mathcal{G})$ to $C^{*}(\mathcal{G})$ without requiring $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}(\mathcal{G}^{x}_{x})$ for $\mu$-almost every $x\in\mathcal{G}^{(0)}$. More precisely, we have the following:
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid which is second-countable, and $\tau$ be a tracial state on $C^{*}(\mathcal{G})$ with the associated measure $\mu$. Then there exists a unique $\mu$-measurable field $\{\tau_{x}\}_{x\in\mathcal{G}^{(0)}}$, where $\tau_{x}$ is a tracial state on $C^{*}(\mathcal{G}_{x}^{x})$ such that
(4.4) $\tau(a)=\int_{\mathcal{G}^{(0)}}\tau_{x}(\vartheta_{x}(a))\mathrm{d}\mu(x)\quad\text{for}\quad a\in C^{*}(\mathcal{G}).$
@@ 5.1. Transformation groupoids
In this subsection, we apply the results in previous sections to transformation groupoids. If $\Gamma$ is a discrete group acting on a locally compact Hausdorff space $X$, then we form the transformation groupoid $X\rtimes\Gamma$. In this case, the reduced groupoid $C^{*}$-algebra $C^{*}_{r}(X\rtimes\Gamma)$ is *-isomorphic to the reduced crossed product $C_{0}(X)\rtimes_{r}\Gamma$, and the maximal groupoid $C^{*}$-algebra $C^{*}(X\rtimes\Gamma)$ is *-isomorphic to the maximal crossed product $C_{0}(X)\rtimes\Gamma$.
For each $\gamma\in\Gamma$, we use the same notation to denote the associated homeomorphism $\gamma:X\ni x\mapsto\gamma x\in X$. Therefore, a probability Radon measure $\mu$ on $X$ is invariant in the sense in Section 2.3 if and only if $\gamma$ is $\mu$-preserving for each $\gamma\in\Gamma$.
The notion of essential freeness for transformation groupoids can be easily simplified as follows:
@@
Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ with an invariant probability Radon measure $\mu$ on $X$. Then the transformation groupoid $X\rtimes\Gamma$ is essentially free with respect to $\mu$ (in the sense of Definition 3.3) if and only if the action is essentially free with respect to $\mu$ in the sense that $\mu(\mathrm{Fix}(\gamma))=0$ for every $\gamma\neq 1$ in $\Gamma$.
Proof.
The necessity follows from the inner regularity of the Radon measure $\mu$, while the sufficiency follows from the fact that every pre-compact bisection in $X\rtimes\Gamma$ is contained in $X\times F$ for some finite $F\subseteq\Gamma$. ∎
@@
where $f_{i}\in C_{0}(X)$ and $\gamma_{i}\in\Gamma$ for $i=1,\dotsc,n$.
@@
Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ with an invariant probability Radon measure $\mu$. Consider the following conditions:
@@
Then (2) $\Rightarrow$ (1).
If $\Gamma$ is countable and $X$ is second-countable, then we actually have (1) $\Leftrightarrow$ (2).
@@
First, we show that “(2) $\Rightarrow$ (1)”. Recall from Lemma 4.2 (see also Remark 4.3) that
@@
holds for any $f\in C_{c}(\mathcal{G})$. By (2), the trivial representation $\tau^{\mathrm{triv}}_{x}$ on $C_{c}(\Gamma_{x})$ extends to a state $\tau^{\mathrm{triv}}_{x,r}$ on $C^{*}_{r}(\Gamma_{x})$ for $\mu$-almost every $x\in X$. By the same argument as in the proof of Proposition 4.9 without requiring the second-countability of $X\rtimes\Gamma$ (see Remark 4.3), the map
@@
Now we show that “(1) $\Rightarrow$ (2)” if $\Gamma$ is countable and $X$ is second-countable. By Proposition 4.6, $\tau^{\mathrm{Fix}}_{\mu}$ extends to a tracial state on $C_{0}(X)\rtimes_{r}\Gamma$ if and only if for $\mu$-almost every $x\in X$, the trivial representation on $C_{c}((X\rtimes\Gamma)_{x}^{x})$ extends to a state on $C^{*}_{e}((X\rtimes\Gamma)_{x}^{x})$. Note that $(X\rtimes\Gamma)_{x}^{x}$ is isomorphic to $\Gamma_{x}$, and it follows from [11, Proposition 2.10] that $\|\cdot\|_{e}=\|\cdot\|_{r}$ on $C_{c}((X\rtimes\Gamma)_{x}^{x})$. It is also known that the trivial representation on $C_{c}(\Gamma_{x})$ extends to a state on $C^{*}_{r}(\Gamma_{x})$ if and only if $\Gamma_{x}$ is an amenable group. Thus, we have concluded the proof. ∎
@@
Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ with an invariant probability Radon measure $\mu$. We consider the following conditions:
@@
Then (1) $\Rightarrow$ (2) and (2) + (3) $\Rightarrow$ (1). In particular, if the action is essentially free then any tracial state $\tau$ on $C_{0}(X)\rtimes_{r}\Gamma$ is canonical.
If additionally $\Gamma$ is countable and $X$ is second-countable, then (1) $\Leftrightarrow$ (2) + (3).
@@
Note that “(2) $\Rightarrow$ (1)” in Corollary 5.5 does not hold in general. Indeed, it was shown in [8, Corollary 1.4] that if the amenable radical of $\Gamma$ (i.e., the largest amenable normal subgroup of $\Gamma$) is trivial (e.g., $\Gamma$ is $C^{*}$-simple), then any tracial state on the reduced crossed product is canonical. Actually, it was shown in [34, Corollary 1.12] that any tracial state on the reduced crossed product is canonical if and only if the action of the amenable radical of $\Gamma$ on $X$ is essentially free. On the other hand, trivial actions by non-trivial groups cannot be essentially free.
@@
Let $\Gamma$ be a discrete group acting trivially on a compact Hausdorff space $X$ with a probability Radon measure $\mu$. Then the following are equivalent:
@@
“(2) $\Leftrightarrow$ (3)”: If $\Gamma$ is amenable, then $X\rtimes\Gamma$ is amenable and hence has the weak containment property. Conversely, if we assume that $X\rtimes\Gamma$ has the weak containment property, then the canonical quotient map $C(X)\otimes_{\text{max}}C^{*}(\Gamma)=C(X)\rtimes\Gamma\to C(X)\rtimes_{r}\Gamma=C(X)\otimes_{\text{min}}C_{r}^{*}(\Gamma)$ is injective, which implies that $C^{*}(\Gamma)\to C^{*}_{r}(\Gamma)$ is injective by a diagram chase. Therefore, $\Gamma$ is amenable. ∎
@@
Let $\Gamma$ be a non-trivial discrete group acting trivially on a compact Hausdorff space $X$ with a probability Radon measure $\mu$. Then at least one of the following conditions fails: